OpenStudy (anonymous):

Need help solving this problem Five years ago, Ed was 2/3 as old as Jeff. Ten years from now, he will be 5/6 as old as Jeff will be. How old is each now?

OpenStudy (anonymous):

let Ed's age be x, and Jeff's be y. (x - 5) = 2/3 * (y - 5) (x + 10) = 5/6 *(y + 10) 3x - 15 = 2y - 10 --> eq. 1 6x + 60 = 5y + 50 --> eq. 2 (2) - (1)*2 --> 90 = y + 70 y = 20, then x = 15